# Yibin Wang ${}^{\text{\textdagger}1}$ &Haizhou Shi ${}^{*\text{\textdagger}1}$ &Ligong Han 12

### v3

@@ A.1 Proof of Main Theorems
${\bm{\Sigma}}_{q}$ $=[{\bm{I}}_{n}\otimes{\bm{B}}]\cdot[\operatorname{diag}(\operatorname{vec}({\bm{\Omega}})^{2})]\cdot[{\bm{I}}_{n}\otimes{\bm{B}}^{\top}].$
@@
Suppose the deterministic matrix ${\bm{B}}=[{\bm{b}}_{1},{\bm{b}}_{2},\cdots,{\bm{b}}_{r}]\in\mathbb{R}^{m\times r}$, random matrix ${\bm{A}}=[{\bm{a}}_{1},{\bm{a}}_{2},\cdot,{\bm{a}}_{r}]^{\top}\in\mathbb{R}^{r\times n}$, with its underlying parameters of mean and standard deviation defined likewise ${\bm{M}}\in\mathbb{R}^{r\times n}$ and ${\bm{\Omega}}\in\mathbb{R}^{r\times n}$. We have ${\bm{W}}={\bm{B}}{\bm{A}}=\sum_{i=1}^{r}{\bm{b}}_{i}\cdot{\bm{a}}_{i}^{\top}$. We then rewrite $\operatorname{vec}({\bm{W}})$ in the form of Kronecker product $\otimes$:
$\operatorname{vec}({\bm{W}})$ $=\operatorname{vec}(\sum_{i=1}^{r}{\bm{b}}_{i}\cdot{\bm{a}}_{i}^{\top})=\sum_{i=1}^{r}({\bm{a}}_{i}\otimes{\bm{b}}_{i})$ (17)
@@
$=\sum_{i=1}^{r}\mathbb{E}_{{\bm{a}}_{i}}[({\bm{a}}_{i}\otimes{\bm{b}}_{i})({\bm{a}}_{i}\otimes{\bm{b}}_{i})^{\top}]$ (20)
$=\sum_{i=1}^{r}\mathbb{E}_{{\bm{a}}_{i}}[({\bm{a}}_{i}{\bm{a}}_{i}^{\top})\otimes({\bm{b}}_{i}{\bm{b}}_{i}^{\top})]=\sum_{i=1}^{r}\mathbb{E}_{{\bm{a}}_{i}}[({\bm{a}}_{i}{\bm{a}}_{i}^{\top})]\otimes({\bm{b}}_{i}{\bm{b}}_{i}^{\top})$ (21)
@@
$=[{\bm{I}}_{n}\otimes{\bm{B}}]\cdot[\operatorname{diag}(\operatorname{vec}({\bm{\Omega}}^{2}))]\cdot[{\bm{I}}_{n}\otimes{\bm{B}}^{\top}],$ (23)
@@
For the prior distribution, as suggested in Eqn. 3.2, we assume
@@
By default, we assume that the low-rank tall matrix $\widetilde{{\bm{R}}}\in\mathbb{R}^{(mn)\times r^{\prime}}$ has the full column rank $r^{\prime}$. Otherwise if $\operatorname{r}(\widetilde{{\bm{R}}})=r^{\prime\prime}<r^{\prime}$, then we can in effect consider a new matrix component $\widetilde{{\bm{R}}}^{\prime}\in\mathbb{R}^{(mn)\times r^{\prime\prime}}$ that has the same rank as $r^{\prime\prime}$, which satisfies our assumption of full column rank. Therefore, we assume that the result of the SVD decomposition of $\widetilde{{\bm{R}}}$ is given by
@@
where ${\bm{U}}_{R}\in\mathbb{R}^{(mn)\times(mn)}$ and ${\bm{V}}_{R}\in\mathbb{R}^{r^{\prime}\times r^{\prime}}$ are orthogonal, i.e., ${\bm{U}}_{R}{\bm{U}}_{R}^{\top}={\bm{U}}_{R}^{\top}{\bm{U}}_{R}={\bm{I}}_{(mn)}$ and ${\bm{V}}_{R}{\bm{V}}_{R}^{\top}={\bm{V}}_{R}^{\top}{\bm{V}}_{R}={\bm{I}}_{r^{\prime}}$. ${\bm{D}}_{R}$ is a tall matrix where its upper part is diagonal and the lower part is a zero matrix, denoted as ${\bm{D}}_{R}=[{\bm{D}}_{R}^{*},{\bm{O}}]^{\top}=[\operatorname{diag}([d_{R_{1}}>0,d_{R_{2}}>0,\cdots,d_{R_{r^{\prime}}}>0]),{\bm{O}}]^{\top}$.
@@
where we simplify the notation for the covariance matrix ${\bm{\Sigma}}_{q}$ by defining $\widetilde{{\bm{B}}}=[{\bm{I}}_{n}\otimes{\bm{B}}]\in\mathbb{R}^{(mn)\times(mr)}$ and ${\bm{\Sigma}}=\operatorname{diag}(\operatorname{vec}({\bm{\Omega}})^{2})$. Likewise, we have the SVD-decomposed matrices for $\widetilde{{\bm{B}}}$ where they are defined in the same way as Eqn. 28:
@@
First, we calculate the log-determinant part of the KL-divergence. For the log-determinant of the covariance matrix of the prior distribution ${\bm{\Sigma}}_{p}$, by applying SVD decomposition in Eqn. 28, we have
@@
We make two observations when $\lambda\rightarrow 0^{+}$: (i) compare the terms that contain $\log\lambda$ on both sides, to make sure the log-determinant in the divergence term bounded, we have to set $r^{\prime}=mr$; (ii) the last term in Eqn. 38, $\log|{\bm{I}}+\lambda{\bm{V}}_{B}^{\top}{\bm{\Sigma}}^{-1}{\bm{V}}_{B}({\bm{D}}_{B}^{*})^{-2}|=\log|{\bm{I}}|=0$. Therefore, we have
@@
Next we calculate $\operatorname{tr}({\bm{\Sigma}}_{p}^{-1}{\bm{\Sigma}}_{q})$ in Eqn. 26. Following the same assumptions and notations above, we have the inverse of the covariance of the prior distribution as
@@
Substituting ${\bm{U}}_{R},{\bm{D}}_{R},{\bm{V}}_{R}$ back to Eqn. 41 and applying $\lambda\rightarrow 0^{+}$, we have
@@
For the quadratic term in Eqn. 26, the pre-trained weights ${\bm{W}}_{0}$ cancel out, and we can calculate it as
@@
Finally, proof is completed by combining Eqn. 39, Eqn. 48, and Eqn. 54. ∎
@@ A.2 Analysis on BLoB Parameterization
Due to the fact that $\sigma_{q}$ is typically initialized to a small value close to 0 to ensure stable optimization of the likelihood cost term (e.g., $31e-$ 3), and in order to ensure that the model obtains good uncertainty, $\sigma_{p}$ is usually set to a larger value close to 1 (e.g., $11e-$ 1). In this case, the derivative of $\rho$ in Eqn. 62 is almost always a constant $-1$, which, based on our previous analysis, leads to slow convergence for large $\sigma_{p}$ values when $\sigma_{q}$ is small.
@@
Under the same initialization conditions, the derivative in Eqn. 63 is approximately of the order of $\rho^{-1}$, leading to rapid convergence towards larger $\sigma_{p}$ values when $\sigma_{q}$ is small. Building on this, we use SGD without momentum to optimize the complexity loss term, thereby achieving the natural convergence of $\sigma_{q}$.
Visual Demonstration. To visually demonstrate the differences in the convergence of KL divergence during training with these two parameterizations, we set $\sigma_{p}=1$ and employ gradient descent to optimize the KL divergence. As introduced in B.1, in practical mini-batch gradient descent, the KL divergence is weighted by $\nicefrac{{1}}{{\text{\#mini-batches}}}$. Therefore, assuming there are 100 mini-batches, the learning rate is set to 0.01, which translates to an actual learning rate of $41e-$ 4 for $\rho$. We initialize $\sigma_{q}=0.01$ for both parameterizations. The growth of $\sigma_{q}$ during the KL training, for $\sigma_{q}=\text{log}(1+e^{\rho})$ and $\sigma_{q}=\rho^{2}$ is shown in Fig. 2. In the same setting, the softplus parameterization takes nearly 100,000 iterations to converge, while the square parameterization takes only about 5,000 iterations.
Figure 2: The growth curve of $\sigma_{q}=\text{log}(1+e^{\rho})$ and $\sigma_{q}=\rho^{2}$ during the optimization of KL divergence (without data likelihood). The number of training steps (5000) is marked with the red line.
@@ B.1 Implementation of BLoB
The pseudo-rescaling does not affect the likelihood cost in practical mini-batch gradient descent. In fact, it only changes the warm-up period in KL reweighting from $M$ to $\nicefrac{{L^{\ast}}}{{\text{batch size}}}$, thereby facilitating more consistent optimization of the complexity cost across datasets of different sizes.

### v5

@@ 6 Conclusion
7 Limitations
The main limitations of our proposed BLoB method are: (i) BLoB is confined to fine-tuning scenarios and is not applicable to training-free tasks, such as direct uncertainty estimation during inference [63]. (ii) As a typical mean-field variational inference method, BLoB requires multiple sampling iterations during inference, which challenges stable and efficient deployment. (iii) While BLoB’s effectiveness has been empirically demonstrated for downstream classification tasks, its application to generation tasks requires further investigation.
Acknowledgement
The authors thank the reviewers/AC for the constructive comments to improve the paper. We thank Sanket Jantre for identifying improvements to our proof and for other valuable discussions. HS and HW are partially supported by Microsoft Research AI & Society Fellowship, NSF Grant IIS-2127918, NSF CAREER Award IIS-2340125, NIH Grant 1R01CA297832, and the Amazon Faculty Research Award. This research is also supported by NSF National Artificial Intelligence Research Resource (NAIRR) Pilot. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the sponsors.
@@ A.1 Proof of Main Theorems
${\bm{\Sigma}}_{q}$ $=[{\bm{I}}_{n}\otimes{\bm{B}}][\operatorname{diag}(\operatorname{vec}({\bm{\Omega}})^{2})][{\bm{I}}_{n}\otimes{\bm{B}}^{\top}].$
@@
Suppose the deterministic matrix ${\bm{B}}=[{\bm{b}}_{1},{\bm{b}}_{2},\cdots,{\bm{b}}_{r}]\in\mathbb{R}^{m\times r}$, random matrix ${\bm{A}}=[{\bm{a}}_{1},{\bm{a}}_{2},\cdots,{\bm{a}}_{r}]^{\top}\in\mathbb{R}^{r\times n}$, with its underlying parameters of mean and standard deviation defined likewise ${\bm{M}}\in\mathbb{R}^{r\times n}$ and ${\bm{\Omega}}\in\mathbb{R}^{r\times n}$. We have ${\bm{W}}={\bm{B}}{\bm{A}}=\sum_{i=1}^{r}{\bm{b}}_{i}{\bm{a}}_{i}^{\top}$. We then rewrite $\operatorname{vec}({\bm{W}})$ in the form of Kronecker product $\otimes$:
$\operatorname{vec}({\bm{W}})$ $=\operatorname{vec}(\sum_{i=1}^{r}{\bm{b}}_{i}{\bm{a}}_{i}^{\top})=\sum_{i=1}^{r}({\bm{a}}_{i}\otimes{\bm{b}}_{i})$ (18)
@@
$=\sum_{i=1}^{r}\left\{\mathbb{E}_{{\bm{a}}_{i}}[({\bm{a}}_{i}\otimes{\bm{b}}_{i})({\bm{a}}_{i}\otimes{\bm{b}}_{i})^{\top}]-\mathbb{E}_{{\bm{a}}_{i}}[({\bm{a}}_{i}\otimes{\bm{b}}_{i})]\mathbb{E}_{{\bm{a}}_{i}}[({\bm{a}}_{i}\otimes{\bm{b}}_{i})^{\top}]\right\}$ (21)
$=\sum_{i=1}^{r}\left\{\mathbb{E}_{{\bm{a}}_{i}}[({\bm{a}}_{i}{\bm{a}}_{i}^{\top})]\otimes({\bm{b}}_{i}{\bm{b}}_{i}^{\top})-(\mathbb{E}_{{\bm{a}}_{i}}[{\bm{a}}_{i}]\mathbb{E}_{{\bm{a}}_{i}}[{\bm{a}}_{i}]^{\top})\otimes({\bm{b}}_{i}{\bm{b}}_{i}^{\top})\right\}$ (22)
@@
$=[{\bm{I}}_{n}\otimes{\bm{B}}][\operatorname{diag}(\operatorname{vec}({\bm{\Omega}}^{2}))][{\bm{I}}_{n}\otimes{\bm{B}}^{\top}],$ (24)
@@
For the prior distribution, following the alternative form of a degenerate Gaussian [83], as suggested in Eqn. 3.2, we assume
@@
By default, we assume that the low-rank tall matrix $\widetilde{{\bm{R}}}\in\mathbb{R}^{(mn)\times r^{\prime}}$ has the full column rank $r^{\prime}$. Otherwise if $\operatorname{r}(\widetilde{{\bm{R}}})=r^{\prime\prime}<r^{\prime}$, then we can in effect consider a new matrix component $\widetilde{{\bm{R}}}^{\prime}\in\mathbb{R}^{(mn)\times r^{\prime\prime}}$ that has the same rank as $r^{\prime\prime}$, which satisfies our assumption of full column rank. Therefore, we have the SVD decomposition of $\widetilde{{\bm{R}}}$ is given by
@@
where ${\bm{U}}_{R}\in\mathbb{R}^{(mn)\times(mn)}$ and ${\bm{V}}_{R}\in\mathbb{R}^{r^{\prime}\times r^{\prime}}$ are orthonormal, i.e., ${\bm{U}}_{R}{\bm{U}}_{R}^{\top}={\bm{U}}_{R}^{\top}{\bm{U}}_{R}={\bm{I}}_{(mn)}$ and ${\bm{V}}_{R}{\bm{V}}_{R}^{\top}={\bm{V}}_{R}^{\top}{\bm{V}}_{R}={\bm{I}}_{r^{\prime}}$. ${\bm{D}}_{R}$ is a tall matrix where its upper part is diagonal and the lower part is a zero matrix, denoted as ${\bm{D}}_{R}=[{\bm{D}}_{R}^{*},{\bm{O}}]^{\top}=[\operatorname{diag}([d_{R_{1}}>0,d_{R_{2}}>0,\cdots,d_{R_{r^{\prime}}}>0]),{\bm{O}}]^{\top}$.
@@
where we simplify the notation for the covariance matrix ${\bm{\Sigma}}_{q}$ by defining $\widetilde{{\bm{B}}}=[{\bm{I}}_{n}\otimes{\bm{B}}]\in\mathbb{R}^{(mn)\times(mr)}$ and ${\bm{\Sigma}}=\operatorname{diag}(\operatorname{vec}({\bm{\Omega}})^{2})$. Likewise, we have the SVD-decomposed matrices for $\widetilde{{\bm{B}}}$ where they are defined in the same way as Eqn. 29:
@@
First, we calculate the log-determinant part of the KL-divergence. For the log-determinant of the covariance matrix of the prior distribution ${\bm{\Sigma}}_{p}$, by applying SVD decomposition in Eqn. 29, we have
@@
We make two observations when $\lambda\rightarrow 0^{+}$: (i) compare the terms that contain $\log\lambda$ on both sides, to make sure the log-determinant in the divergence term bounded, we have to set $r^{\prime}=mr$; (ii) the last term in Eqn. 39, $\log|{\bm{I}}+\lambda{\bm{V}}_{B}^{\top}{\bm{\Sigma}}^{-1}{\bm{V}}_{B}({\bm{D}}_{B}^{*})^{-2}|=\log|{\bm{I}}|=0$. Therefore, we have
@@
Next we calculate $\operatorname{tr}({\bm{\Sigma}}_{p}^{-1}{\bm{\Sigma}}_{q})$ in Eqn. 27. Following the same assumptions and notations above, we have the inverse of the covariance of the prior distribution as
@@
Substituting ${\bm{U}}_{R},{\bm{D}}_{R},{\bm{V}}_{R}$ back to Eqn. 42 and applying $\lambda\rightarrow 0^{+}$, we have
@@
For the quadratic term in Eqn. 27, the pre-trained weights ${\bm{W}}_{0}$ cancel out, and we can calculate it as
@@
Finally, proof is completed by combining Eqn. 40, Eqn. 49, and Eqn. 55. ∎
@@ A.2 Analysis on BLoB Parameterization
Due to the fact that $\sigma_{q}$ is typically initialized to a small value close to 0 to ensure stable optimization of the likelihood cost term (e.g., $31e-$ 3), and in order to ensure that the model obtains good uncertainty, $\sigma_{p}$ is usually set to a larger value close to 1 (e.g., $11e-$ 1). In this case, the derivative of $\rho$ in Eqn. 63 is almost always a constant $-1$, which, based on our previous analysis, leads to slow convergence for large $\sigma_{p}$ values when $\sigma_{q}$ is small.
@@
Under the same initialization conditions, the derivative in Eqn. 64 is approximately of the order of $\rho^{-1}$, leading to rapid convergence towards larger $\sigma_{p}$ values when $\sigma_{q}$ is small. Building on this, we use SGD without momentum to optimize the complexity loss term, thereby achieving the natural convergence of $\sigma_{q}$.
Visualization. To visually demonstrate the differences in the convergence of KL divergence during training with these two parameterizations, we set $\sigma_{p}=1$ and employ gradient descent to optimize the KL divergence. As introduced in B.1, in practical mini-batch gradient descent, the KL divergence is weighted by $\nicefrac{{1}}{{\text{\#mini-batches}}}$. Therefore, assuming there are 100 mini-batches, the learning rate is set to 0.01, which translates to an actual learning rate of $41e-$ 4 for $\rho$. We initialize $\sigma_{q}=0.01$ for both parameterizations. The growth of $\sigma_{q}$ during the KL training, for $\sigma_{q}=\text{log}(1+e^{\rho})$ and $\sigma_{q}=\rho^{2}$ is shown in Fig. 2. In the same setting, the softplus parameterization takes nearly 100,000 gradient steps to converge, while the square parameterization takes only about 5,000 gradient steps. This modification makes it suitable for our fine-tuning setup, where the number of gradient steps is relatively small.
Figure 2: The growth curve of $\sigma_{q}=\text{log}(1+e^{\rho})$ and $\sigma_{q}=\rho^{2}$ during the optimization of KL divergence (without data likelihood). The number of gradient steps (5000) is marked with the red line.
@@ B.1 Implementation of BLoB
Pseudo-rescaling does not affect the likelihood cost in practical mini-batch gradient descent. It modifies the warm-up period in KL reweighting from $M$ to $\nicefrac{{L^{\ast}}}{{\text{batch size}}}$, enabling more consistent optimization of the complexity cost across datasets of different sizes. Pseudo-rescaling is specifically designed for training procedures that use a fixed number of iterations across different datasets. However, for training procedures with a fixed number of epochs across different datasets, pseudo-rescaling can be omitted since the complexity cost converges at an identical rate across all datasets.
In Table 5, we present the results of BLoB without pseudo-rescaling after 5 epochs of LoRA fine-tuning, alongside other baseline methods. In this setting, we increase the maximum learning rate from 1e-4 to 2e-4. Fine-tuning for 5 epochs reduces the number of gradient steps required for methods like MLE on smaller datasets such as WG-S, resulting in less overfitting (as evidenced by lower ECE and NLL) while maintaining comparable accuracy to fine-tuning with 5000 gradient steps. BLoB continues to demonstrate superior uncertainty estimation performance across nearly all datasets.
@@
Table 5: Comparison of BLoB without Pseudo-rescaling against baseline methods applied to LoRA on Llama2-7B pre-trained weights, where Accuracy (ACC) and Expected Calibration Error (ECE) are reported in percentages. The evaluation is done across six common-sense reasoning tasks with a shared hyper-parameter setting after 5 epochs. We use $N$ to represent the number of samples during inference in BLoB. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface and underlining denote the best and the second-best performance, respectively.
Metric Method Datasets
WG-S [82] ARC-C [18] ARC-E [18] WG-M [82] OBQA [64] BoolQ [17]
ACC ($\uparrow$) MLE 68.56$\pm$0.52 71.39$\pm$0.52 86.49$\pm$0.42 76.81$\pm$0.17 81.32$\pm$0.80 88.09$\pm$0.21
MAP 68.96$\pm$0.65 69.37$\pm$0.57 86.27$\pm$0.25 77.35$\pm$1.11 82.33$\pm$0.41 87.95$\pm$0.08
MCD [29] 69.15$\pm$1.23 68.81$\pm$0.16 87.50$\pm$0.63 76.64$\pm$0.89 81.65$\pm$0.75 88.04$\pm$0.59
ENS [50, 8, 103] 67.88$\pm$0.79 69.14$\pm$1.11 87.73$\pm$0.17 76.19$\pm$1.11 82.06$\pm$0.57 88.10$\pm$0.17
LAP [116] 69.01$\pm$0.49 58.22$\pm$6.85 66.49$\pm$9.56 76.53$\pm$1.64 82.66$\pm$0.49 87.96$\pm$0.15
BLoB (N=0) 68.08$\pm$1.18 70.42$\pm$2.23 86.85$\pm$0.30 74.89$\pm$1.19 82.99$\pm$0.69 88.60$\pm$0.35
BLoB (N=10) 66.69$\pm$1.12 70.00$\pm$1.80 86.37$\pm$0.47 75.17$\pm$0.76 82.92$\pm$0.74 88.03$\pm$0.27
ECE ($\downarrow$) MLE 19.56$\pm$0.56 22.49$\pm$1.11 11.33$\pm$0.51 18.37$\pm$0.52 15.60$\pm$0.32 9.65$\pm$0.07
MAP 19.03$\pm$1.86 24.15$\pm$0.90 11.50$\pm$0.35 17.29$\pm$0.64 14.33$\pm$0.78 9.88$\pm$0.13
MCD [29] 18.69$\pm$1.37 23.94$\pm$1.02 10.34$\pm$0.36 16.78$\pm$0.41 14.51$\pm$0.99 9.25$\pm$0.58
ENS [50, 8, 103] 14.86$\pm$3.11 21.86$\pm$0.94 9.36$\pm$0.36 9.21$\pm$2.16 12.30$\pm$0.90 7.65$\pm$0.02
LAP [116] 4.16$\pm$0.58 19.27$\pm$4.13 27.21$\pm$4.08 7.80$\pm$0.98 14.73$\pm$1.78 5.76$\pm$1.59
BLoB (N=0) 14.93$\pm$0.97 20.51$\pm$0.72 9.63$\pm$0.38 13.31$\pm$0.45 9.95$\pm$0.98 7.13$\pm$0.20
BLoB (N=10) 10.43$\pm$0.71 11.93$\pm$0.50 5.38$\pm$0.38 6.82$\pm$1.29 3.67$\pm$0.66 3.73$\pm$0.48
NLL ($\downarrow$) MLE 0.82$\pm$0.03 1.48$\pm$0.10 0.86$\pm$0.07 0.88$\pm$0.01 1.15$\pm$0.06 0.67$\pm$0.01
MAP 0.83$\pm$0.08 1.53$\pm$0.10 0.79$\pm$0.02 0.85$\pm$0.01 0.97$\pm$0.05 0.74$\pm$0.02
MCD [29] 0.80$\pm$0.04 1.53$\pm$0.03 0.77$\pm$0.06 0.80$\pm$0.01 0.97$\pm$0.09 0.65$\pm$0.05
ENS [50, 8, 103] 0.72$\pm$0.06 1.30$\pm$0.05 0.58$\pm$0.00 0.56$\pm$0.03 0.74$\pm$0.04 0.46$\pm$0.01
LAP [116] 0.61$\pm$0.00 1.46$\pm$0.02 1.41$\pm$0.02 0.55$\pm$0.00 1.04$\pm$0.01 0.47$\pm$0.01
BLoB (N=0) 0.70$\pm$0.02 1.07$\pm$0.03 0.56$\pm$0.01 0.63$\pm$0.00 0.57$\pm$0.03 0.39$\pm$0.01
BLoB (N=10) 0.65$\pm$0.01 0.84$\pm$0.01 0.44$\pm$0.00 0.54$\pm$0.01 0.46$\pm$0.01 0.31$\pm$0.00